# Rings in which one-sided strongly $\pi$-regular elements are strongly $\pi$-regular

### v1

@@ Abstract
In 1977, Hartwig and Luh asked if $a$ is an element of a Dedekind-finite ring $S$, then does $aS=a^{2}S$ imply $Sa=Sa^{2}$. This question was answered negatively by Dittmer, Khurana, and Nielsen in 2014. On the other hand, Dittmer et al. proved that the question of Hartwig and Luh has a positive answer for Dedekind-finite exchange rings. We explore the question of Hartwig and Luh for various other classes of Dedekind-finite rings. We will also prove that the condition in question is not left-right symmetric.
@@ 1 Introduction
Throughout this paper, all rings are assumed to be associative, but not necessarily commutative. The ring $R$ is not assumed to have an identity. Whenever $R$ possesses an identity, which will be denoted by $1$. Recall that a ring $R$ is called strongly regular if for every $a\in R$ there exists an $x\in R$ depending on the element $a$ such that $a=a^{2}x$. A ring $R$ is said to be strongly $\pi$-regular if for every $a\in R$ there exist $x\in R$ and $n\in\mathbb{N}$ (both depending on the element $a$) such that $a^{n}=a^{n+1}x$. If any ring $R$ has an identity element, then $R$ is called strongly $\pi$-regular if for every $a\in R$ the descending chain $aR\supseteq a^{2}R\supseteq\ldots$ of the principal right ideals terminates. Similarly, any ring $R$ with identity element is called strongly regular if for every $a\in R$, $aR=a^{2}R$. There is no need to study the left and right versions of these definitions separately, as these are left-right symmetric in view of the seminal work of Arens and Kaplansky [1] and Dischinger [4].
Recall that an element $a$ of a ring $R$ is said to be right (respectively, left) strongly $\pi$-regular if $a^{n}=a^{n+1}x$ (respectively, $a^{n}=xa^{n+1}$) for some $x\in R$ and $n\in\mathbb{N}$. If $n=1$, then $a$ is called a right (left) strongly regular element. An element which is both left and right strongly $\pi$-regular (respectively, strongly regular) is called a strongly $\pi$-regular (respectively, strongly regular) element. Arens and Kaplansky [1] introduced strongly regular rings as a generalization of regular rings. Subsequently, Azumaya introduced strongly $\pi$-regular rings as a generalization of $\pi$-regular rings. In [7], Kaplansky asked the following question
@@
In 1954, Azumaya [2, Theorem 3] proved that if $a\in R$ is strongly $\pi$-regular, then there exists an element $z$ commuting with $a$ and $n\in\mathbb{N}$ such that $a^{n}=a^{n+1}z$. This implies that $a^{n}=a^{2n}z^{n}=a^{n}z^{n}a^{n}$ and so $a^{n}z^{n}$ is an idempotent. Thus, a strongly $\pi$-regular element enables one to construct idempotents as desired by Kaplansky in the question mentioned above. But when Kaplansky raised this question, it was not known that if all elements of a ring are right strongly $\pi$-regular, then all its elements are also left strongly $\pi$-regular. This result was proved about $25$ years later by Dischinger [4] in 1976. This result of Dischinger [4] leads to a very satisfactory answer to the question raised by Kaplansky. However, this result of Dischinger does not hold locally, and a right strongly $\pi$-regular element may not be left strongly $\pi$-regular. For instance, if $R$ is not Dedekind finite and has two elements a,ba,\,b such that $ab=1$ but $ba\neq 1$, then $a$ is right strongly $\pi$-regular but not left strongly $\pi$-regular. On the other hand, Azumaya [2, Theorem 4] proved that if $R$ has a bounded index of nilpotence, then every one-sided strongly $\pi$-regular element is strongly $\pi$-regular. Azumaya [2, Lemma 3] also proved that if $a\in R$ is strongly $\pi$-regular with $a^{n}=a^{n+1}x$ and $a^{m}=ya^{m+1}$ for some $x,\,y\in R$ and $n,\,m\in\mathbb{N}$, then $a^{m}=a^{m+1}x$ and $a^{n}=ya^{n+1}$ also. So the following result follows.
@@
We call a ring $R$ to be right Dischinger if it satisfies any of the equivalent conditions of Proposition 1.1. Left Dischinger rings are defined similarly. A ring which is both left and right Dischinger will be simply called a Dischinger ring. In other words, Dischinger rings are those in which Dischinger’s result also holds locally.
It is clear that if $R$ is a one-sided Dischinger ring with identity, then $R$ must be Dedekind-finite. Hartwig and Luh [6, Page 94] in 1977 asked whether a Dedekind-finite ring is right Dischinger. A strong negative answer to this question was given by Dittmer, Khurana, and Pace in [5, Theorem 2.3] by constructing an example of a ring with only trivial idempotents that is not right Dischinger. To be specific, it was proved in [5, Theorem 2.3] that the ring
@@
It is well-known that for any two elements $a$ and $b$ in ring $R$ such that $ab=1$ but $ba\neq 1$, then $E_{ij}=b^{i}(1-ba)a^{j}$, where $i,j\in\mathbb{N}$ is an infinite set of matrix units. In view of this the rings in which nilpotent elements satisfy certain properties like form an ideal or a subring are always Dedekind-finite. Azumaya in [2, Theorem 4] already proved that if $R$ has a bounded index of nilpotence, then every right strongly regular element is strongly regular. It is quite interesting to see under which conditions on the set of nilpotent elements one-sided strongly $\pi$-regular elements becomes strongly $\pi$-regular. Also, almost a decade ago, through private communication, Pace P. Nielsen asked one of the authors whether, in NI rings or more generally, in rings in which the set of nilpotent elements is closed under addition or is right linearly McCoy, Does Dischinger’s result holds locally? Also, Azumaya (see [2, Theorem 2]) proved that a right Noetherian ring is Dischinger. In light of these considerations, we may pose this question within the following well-known classes of Dedekind-finite rings: weakly semicommutative, one-sided duo, one-sided Goldie, one-sided Noetherian, NR, NI, UU, one-sided linearly McCoy rings, etc. We will answer all these questions in this paper. We also generalize the result of Azumaya [2, Theorem 4] that rings with bounded index of nilpotence are Dischinger in Theorem 2.10. We also prove that any ring in which every nil subring is nilpotent is a Dischinger ring. Furthermore, we characterize the left and right zero-divisors, units, and nilpotent elements in the ring appearing in (1.3). We additionally show that the Dischinger property is not symmetric i.e., a left Dischinger ring need not be a right Dischinger ring. Finally, we raise several interesting questions for further investigation.
@@ 2 Some classes of right Dischinger rings
Let $a$ be an element in the ring $R$ such that $a=a^{2}x$ and $a$ is left strongly $\pi$-regular, then $a$ is strongly regular.
Dittmer et al. [5, Lemma 3.3] proved that if $a$ is a regular element in a Dedekind-finite ring $R$ and $aR=a^{2}R$, then $Ra=Ra^{2}$. Using Azumaya’s result [2, Lemma 3], we can extend this result without additional effort as follows.
@@
Let $a$ be an element of a Dedekind-finite ring $R$ such that $a=a^{2}x$ for some $x\in R$. If $a$ is $\pi$-regular, then $Ra=Ra^{2}$.
Suppose $a^{n}$ is regular, that is, for some $x\in R$, $a^{n}=a^{n}ya^{n}$. As $a=a^{2}x$ then $a^{n}=a^{n+1}x=a^{2n}x^{n}$. By [5, Lemma 3.3], $a^{n}$ is strongly regular, that is, $a$ is strongly $\pi$-regular. By Lemma 2.1, $a$ is strongly regular. ∎
The proof of [5, Theorem 3.4] yields the following useful result.
@@
Let $a$ be an element of a Dedekind-finite ring $R$. If there exists an idempotent $e\in R$ such that $aR+eR=R$ and $r_{R}(a)+(1-e)R=R$, then $a$ is unit-regular.
Recall that an element $a\in R$ is called suitable if there exists an idempotent $e\in aR$ such that $1-e\in(1-a)R$. It was proved in [9] that being suitable is a left-right symmetric property. An element $a\in R$ is called right exchange if for any right ideal $I$ of $R$ with $aR+I=R$, there exists an idempotent $e\in aR$ such that $1-e\in I$. An element is called $\pi$-regular if its some power is (von Neumann) regular. Nicholson [11, Proposition 1.6] proved that a regular element is suitable. Using the same idea, it is easy to prove that $\pi$-regular elements are also suitable.
@@
Let $a$ be an element of a Dedekind-finite ring $R$ such that $a=a^{2}x$ for some $x\in R$. Then the following are equivalent.
(1). $Ra=Ra^{2}$.
(2). $ax$ is a suitable element.
(3). $ax$ is an exchange element.
(4). $ax$ is a $\pi$-regular.
(5). $ax$ is a regular element.
(6). $ax$ is an idempotent.
(7). $axa=a$.
@@
The set of all nilpotent elements of a ring $R$ is denoted by $\rm{N}(R)$. A ring $R$ is called weakly semicommutative if for any $a,\,b\in R$ with $ab=0$ implies $aRb\subseteq\rm{N}(R)$.
@@
Let $R$ be a weakly semicommutative ring. It is enough to show that $R$ is right Dischinger. Suppose $a=a^{2}x$ for some $a,\,x\in R$. As a(1ax)=0a(1-ax)=0, so $ax(1-ax)\in\rm{N}(R)$. This implies that $ax$ is strongly $\pi$-regular. So in view of Proposition 2.4, it is enough to show that $R$ is Dedekind-finite. Let $xy=1$ for some $x,\,y\in R$ then x(1yx)=0x(1-yx)=0. Since $R$ is weakly semicommutative, xy(1yx)=1yxxy(1-yx)=1-yx is nilpotent. Also, 1yx1-yx is an idempotent. This can happen only if 1yx=01-yx=0, that is, $yx=1$. ∎
Let $R$ be a ring with identity in which $a\in\rm{N}(R)$ and $r\in R$ imply that $ar\in\rm{N}(R)$. Then $R$ is weakly semicommutative. Indeed suppose $ab=0$ for some $a,\,b\in R$. Then $ba\in\rm{N}(R)$ and so $bar\in\rm{N}(R)$ for every $r\in R$. This implies that $arb\in\rm{N}(R)$ for every $r\in R$. We have repeatedly used the fact that $xy\in\rm{N}(R)$ iff $yx\in\rm{N}(R)$.
@@
If $R$ is left duo (i.e., $Ra$ is a right ideal for any $a\in R$) and $a\in\rm{N}(R)$, it is easy to see that $ra\in\rm{N}(R)$ for any $r\in R$. Indeed if $a^{n}=0$, then as $aR\subseteq Ra$, so $(ar)^{n}\in Ra^{n}R=0$. So by Corollary 2.6, a one-sided duo ring is Dischinger. A ring $R$ is said to be weakly left duo if for every $a\in R$, there exists an $n\in\mathbb{N}$ such that the left ideal $Ra^{n}$ is a two-sided ideal. Next, we establish another sufficient condition (may not necessary) for a Dedekind-finite ring to be right Dischinger.
@@
As $a^{n}\in Ra^{n}$ and $Ra^{n}$ is an ideal, this implies $a^{n}x^{n-1}=a\in Ra^{n}$. If $n>1$, there is nothing to prove. Suppose $n=1$, this implies $Ra$ is an ideal, so $ax=ya$ for some $y\in R$. Now, the left multiplication with $a$ gives $a=aya$. As $R$ is Dedekind-finite, by [5, Lemma 3.3], $a$ is strongly regular. ∎
@@
It is enough to show that $R$ is Dedekind-finite. Let $xy=1$ for some $x,\,y\in R$. As $yx$ is an idempotent, so for every $n\in\mathbb{N}$ we have $R(yx)^{n}=Ryx$ is an ideal of $R$. In particular, $yxy=y\in Ryx$, implying that $R=Ry\subseteq Ryx\subseteq Rx$. Thus $Rx=R$ and so $R$ is Dedekind finite. ∎
Further, if $R$ is an abelian ring, then the converse of Theorem 2.7 is also true.
@@
Since $a$ is strongly regular, we have $Ra=Re$ for some idempotent $e\in R$. As $R$ is an abelian ring, $Ra$ is an ideal. Hence, for any $r\in R$, $a^{n}r\in Ra^{n}$ for any $n\in\mathbb{N}$. ∎
However, it is not necessary that for a strongly regular element $a\in R$, $Ra^{n}$ is an ideal for some $n\in\mathbb{N}$. For example, if $R=\mathbb{M}_{2}(\mathbb{Q})$. Then $a=E_{11}$ is a strongly regular element in $R$, but $Ra^{n}=Ra$ is not an ideal for any $n\in\mathbb{N}.$
@@
Any ring with identity satisfying the ascending chain condition on left annihilators is left Dischinger. In particular, left Noetherian rings are left Dischinger.
It is natural to ask whether a ring satisfying the ascending chain condition on left annihilators is right Dischinger. In Section 3, we will show that this may not be. Moreover, we prove that with some additional condition, a ring satisfying the condition in Theorem 2.11 becomes right Dischinger (see Corollary 2.14). For this, we need the following lemma. We will tacitly use the fact that if $ab$ is nilpotent, then so is $ba$ for any $a,\,b\in R$.
@@
Let $a$ be an element of a ring $R$ and $\rm{N}^{r}_{R}(a)$ be the subset of nilpotent elements such that for every $r\in\rm{N}^{r}_{R}(a)$ there exist elements $r_{i}\in R$, $i\in\mathbb{N}$ such that $r=r_{1}a$, $ar_{i}=r_{i+1}a$ and $ar_{k}=0$ for some $k\in\mathbb{N}$. Then $\rm{N}^{r}_{R}(a)$ is a nil subring of $R$.
It is easy to see that $\rm{N}^{r}_{R}(a)$ is closed under addition. Let $r,\,s\in\rm{N}^{r}_{R}(a)$. Then there exist $r_{i},\,s_{i}\in R$ for $i\in\mathbb{N}$ such that $r=r_{1}a$, $ar_{i}=r_{i+1}a$, $s=s_{1}a$, and $as_{i}=s_{i+1}a$ for all $i\in\mathbb{N}$. Then $ar_{i}as_{i}=r_{i+1}as_{i+1}a$ for all $i$. Let $k$ be the least positive integer such that one of the $ar_{k}$ or $as_{k}$ is zero. Then $ar_{k}as_{k}=0$ implying that $r_{1}as_{1}a\in\rm{N}^{r}_{R}(a)$. ∎
Analogously, we can define the left version of $\rm{N}^{r}_{R}(a)$ for any element $a\in R$ and denote it by $\rm{N}^{\ell}_{R}(a)$.
@@
A ring $R$ is called left Goldie if it satisfies the ascending chain condition on left annihilators and $_{R}R$ has finite uniform dimension. In [10], Lanski proved that every nil subring of a left Goldie ring is nilpotent. In view of this and Theorem 2.13, we have the following result.
@@
A ring $R$ in which every nil subring is nilpotent is Dischinger. In particular, a left Goldie ring is a Dischinger.
As a left Noetherian ring is left Goldie, so one-sided Noetherian rings are Dischinger.

### v2

@@ Abstract
In 1977, Hartwig and Luh asked: If $a$ is an element in a Dedekind-finite ring $S$, then does $aS=a^{2}S$ imply $Sa=Sa^{2}$? This question was answered negatively by Dittmer, Khurana, and Nielsen in 2014. On the other hand, Dittmer et al. proved that the question of Hartwig and Luh has a positive answer for Dedekind-finite exchange rings. We explore the question of Hartwig and Luh for various other classes of Dedekind-finite rings. We will also prove that the condition in question is not left-right symmetric.
@@ 1 Introduction
Kaplansky asked the following question in his 1950 paper [7]:
@@
Following the literature, an element $a$ in a ring $R$ is right strongly $\pi$-regular, if there exists some $x\in R$ such that $a^{n}=a^{n+1}x$ for some integer $n\geq 1$. Kaplansky was interested in this condition because in rings with identity, this definition is equivalent to termination of the descending chain of principal right ideals $aR\supseteq a^{2}R\supseteq a^{3}R\supseteq\ldots$. Similarly, left strongly $\pi$-regular elements are defined. Moreover, if an element $a$ in a ring $R$ is both left and right strongly $\pi$-regular, then $a$ is strongly $\pi$-regular.
On the other hand if $a=a^{2}x$ for $a,\,x\in R$, then $a$ is said to be right strongly regular. The left strongly regular elements are defined similarly. An element $a\in R$ which is both left and right strongly regular is called strongly regular. It is clear that if $a$ is strongly regular, then $R=aR\oplus r_{R}(a)$ implying that every strongly regular element is (von Neumann) regular. It is easy to see that an element is strongly $\pi$-regular iff its some power is strongly regular. As it is common in ring theory to add $\pi$ when talking about conditions that hold for some power, it would have been more appropriate to call strongly $\pi$-regular element as $\pi$-strongly regular.
It is quite easy to see that a strongly $\pi$-regular element leads to an idempotent, as desired by Kaplansky in the above question, as its some power is regular. But the Kaplansky’s question was for rings whose every element is right strongly $\pi$-regular. When Kaplansky raised this question, it was not known that if all elements of a ring are right strongly $\pi$-regular, then all its elements are also left strongly $\pi$-regular. This amazing result was proved about $25$ years later by Dischinger [4] in 1976. This result of Dischinger [4] leads to a very satisfactory answer to the question raised by Kaplansky.
However, this result of Dischinger does not hold locally [^3] The term “locally” is used in the sense of considering the result element-wise, and a right strongly $\pi$-regular element may not be left strongly $\pi$-regular. For instance, if $R$ is not Dedekind finite and has two elements a,ba,\,b such that $ab=1$ but $ba\neq 1$, then $a$ is right strongly $\pi$-regular but not left strongly $\pi$-regular. On the other hand, Azumaya [2, Theorem 4] proved that if $R$ has a bounded index of nilpotence, then every one-sided strongly $\pi$-regular element is strongly $\pi$-regular. Azumaya [2, Lemma 3] also proved that if $a\in R$ is strongly $\pi$-regular with $a^{n}=a^{n+1}x$ and $a^{m}=ya^{m+1}$ for some $x,\,y\in R$ and $n,\,m\in\mathbb{N}$, then $a^{m}=a^{m+1}x$ and $a^{n}=ya^{n+1}$ also. In view of this, the following result follows.
@@
Suppose $(1)$ holds and $a\in R$ such that $a=a^{2}x$, then by assumption, $a$ is strongly $\pi$-regular element i.e., there exists an element $y\in R$ such that $a^{n}=ya^{n+1}$ for some integer $n\geq 1$. By [2, Lemma 3], we have $a=ya^{2}$.
Next, suppose $(2)$ holds and assume that $a\in R$ such that $a^{n}=a^{n+1}x$ for some integer $n\geq 1$ and $x\in R$. Since $a^{n}=a^{n+1}x=a^{2n}x^{n}$, thus by assumption, $a^{n}$ is strongly regular. This implies that there exists an $y\in R$ such that $a^{n}=ya^{2n}$ for some integer $n\geq 1$. Therefore, $a$ is strongly $\pi$-regular. ∎
@@
We call a ring $R$ to be right Dischinger if it satisfies either of the equivalent conditions of Proposition 1.1. Left Dischinger rings are defined similarly. A ring which is both left and right Dischinger will be simply called a Dischinger ring. In other words, Dischinger rings are those in which Dischinger’s result also holds locally.
It is well known that Drazin-invertible elements are precisely the strongly $\pi$-regular elements. Thus in a right Dischinger ring every right strongly $\pi$-regular element is Drazin invertible. Note that the definition of one-sided Dischinger rings is still valid for rings without identity. Throughout this paper, all rings are assumed to be associative but may not have an identity. However, if $R$ possesses an identity, then it will be denoted by $1$. It is clear that if $R$ is a one-sided Dischinger ring with identity, then $R$ must be Dedekind-finite. Hartwig and Luh [6, Page 94] in 1977 asked whether a Dedekind-finite ring is right Dischinger. A strong negative answer to this question was given by Dittmer, Khurana, and Pace in [5, Theorem 2.3] by constructing an example of a ring with only trivial idempotents that is not right Dischinger. To be specific, it was proved in [5, Theorem 2.3] that the ring
@@
It is well-known that if in a ring $R$ there exist $a$ and $b$ such that $ab=1$ but $ba\neq 1$, then $E_{ij}=b^{i}(1-ba)a^{j}$, where $i,j\in\mathbb{N}$ is an infinite set of matrix units. In view of this, the rings in which nilpotent elements satisfy certain properties, such as, nilpotent elements form an ideal (such rings are called NI) or a subring (such rings are called NR), are always Dedekind-finite. Azumaya in [2, Theorem 4] already proved that if $R$ has a bounded index of nilpotence, then every right strongly regular element is strongly regular. It is quite interesting to see under which conditions on the set of nilpotent elements one-sided strongly $\pi$-regular elements becomes strongly $\pi$-regular. Also, almost a decade ago, through private communication, Nielsen asked one of the authors whether Dischinger’s result holds locally in NI rings or more generally, in rings in which the set of nilpotent elements is closed under addition or ring is right linearly McCoy. Also, Azumaya (see [2, Theorem 2]) proved that a right Noetherian ring is Dischinger. In light of these considerations, we may pose this question in well-known classes of Dedekind-finite rings: weakly semicommutative, one-sided duo, one-sided Goldie, one-sided Noetherian, NI, etc. We also generalize the result of Azumaya [2, Theorem 4] that rings with bounded index of nilpotence are Dischinger in Theorem 2.10. Also, by characterizing units, nilpotents, zero-divisors, etc. in the ring appearing in (1.3), we can show that a large number of properties do not yield the Dischinger condition, such as, NR, UU (rings in which units are unipotent), one-sided linearly McCoy, etc. Also, we prove that rings of the form (1.3) helps us to answer that Dischinger condition is not left-right symmetric i.e., one-sided Dischinger ring may not be Dischinger. In the end, we also raise several interesting questions for further investigation.
@@ 2 Some classes of right Dischinger rings
Let $a$ be an element in a ring $R$ such that $a=a^{2}x$ and $a$ is left strongly $\pi$-regular, then $a$ is strongly regular.
Dittmer et al. [5, Lemma 3.3] proved that if $a$ is a regular element in a Dedekind-finite ring $R$ and $aR=a^{2}R$, then $Ra=Ra^{2}$. Using Azumaya’s result [2, Lemma 3], we can extend this result easily as follows.
@@
Let $a$ be an element in a Dedekind-finite ring $R$ such that $a=a^{2}x$ for some $x\in R$. If $a$ is $\pi$-regular, then $Ra=Ra^{2}$.
Suppose $a^{n}$ is regular, that is, for some $y\in R$, $a^{n}=a^{n}ya^{n}$. As $a=a^{2}x$ then $a^{n}=a^{n+1}x=a^{2n}x^{n}$. By [5, Lemma 3.3], $a^{n}$ is strongly regular, that is, $a$ is strongly $\pi$-regular. By Lemma 2.1, $a$ is strongly regular. ∎ The proof of [5, Theorem 3.4] yields the following useful result.
@@
Let $a$ be an element in a Dedekind-finite ring $R$. If there exists an idempotent $e\in R$ such that $aR+eR=R$ and $r_{R}(a)+(1-e)R=R$, then $a$ is unit-regular.
Recall that an element $a\in R$ is called suitable if there exists an idempotent $e\in aR$ such that $1-e\in(1-a)R$. It was proved in [9] that being suitable is a left-right symmetric property. An element $a\in R$ is called right exchange if for any right ideal $I$ of $R$ with $aR+I=R$, there exists an idempotent $e\in aR$ such that $1-e\in I$. Nicholson [11, Proposition 1.6] proved that a regular element is suitable. Using the same idea, it is easy to prove that $\pi$-regular elements are also suitable.
@@
Let $a$ be an element in a Dedekind-finite ring $R$ such that $a=a^{2}x$ for some $x\in R$. Then the following are equivalent.
(1) $Ra=Ra^{2}$.
(2) $ax$ is a suitable element.
(3) $ax$ is an exchange element.
(4) $ax$ is a $\pi$-regular.
(5) $ax$ is a regular element.
(6) $ax$ is an idempotent.
(7) $axa=a$.
@@
The set of all nilpotent elements in a ring $R$ is denoted by $\rm{N}$$(R)$. A ring $R$ is called weakly semicommutative if for any $a,\,b\in R$ with $ab=0$ implies $aRb\subseteq\rm{N}$$(R)$.
@@
Let $R$ be a weakly semicommutative ring. It is enough to show that $R$ is right Dischinger. Suppose $a=a^{2}x$ for some $a,\,x\in R$. As a(1ax)=0a(1-ax)=0, so $ax(1-ax)\in\rm{N}$$(R)$. This implies that $ax$ is strongly $\pi$-regular. So in view of Proposition 2.4, it is enough to show that $R$ is Dedekind-finite. Let $xy=1$ for some $x,\,y\in R$ then x(1yx)=0x(1-yx)=0. Since $R$ is weakly semicommutative, xy(1yx)=1yxxy(1-yx)=1-yx is nilpotent. Also, 1yx1-yx is an idempotent. This can happen only if 1yx=01-yx=0, that is, $yx=1$. ∎ Let $R$ be a ring with identity in which $a\in\rm{N}$$(R)$ and $r\in R$ imply that $ar\in\rm{N}$$(R)$. Then $R$ is weakly semicommutative. Indeed suppose $ab=0$ for some $a,\,b\in R$. Then $ba\in\rm{N}$$(R)$ and so $bar\in\rm{N}$$(R)$ for every $r\in R$. This implies that $arb\in\rm{N}$$(R)$ for every $r\in R$. We have repeatedly used the fact that $xy\in\rm{N}$$(R)$ iff $yx\in\rm{N}$$(R)$.
@@
If $R$ is left duo (i.e., $Ra$ is a right ideal for any $a\in R$) and $a\in\rm{N}$$(R)$, it is easy to see that $ra\in\rm{N}$$(R)$ for any $r\in R$. Indeed if $a^{n}=0$, then as $aR\subseteq Ra$, so $(ar)^{n}\in Ra^{n}R=0$. So by Corollary 2.6, a one-sided duo ring is Dischinger. A ring $R$ is said to be weakly left duo if for every $a\in R$, there exists an $n\in\mathbb{N}$ such that the left ideal $Ra^{n}$ is a two-sided ideal. Next, we establish another sufficient condition (may not be necessary) for a Dedekind-finite ring to be right Dischinger, which will be helpful to prove that weakly left duo rings are right Dischinger.
@@
As $a^{n}\in Ra^{n}$ and $Ra^{n}$ is an ideal, this implies $a^{n}x^{n-1}=a\in Ra^{n}$. If $n>1$, there is nothing to prove. Suppose $n=1$, this implies $Ra$ is an ideal, so $ax=ya$ for some $y\in R$. On left multiplication this with $a$ gives $a=aya$. As $R$ is Dedekind-finite, by [5, Lemma 3.3], $a$ is strongly regular. ∎
@@
In view of the above result, i t is enough to show that $R$ is Dedekind-finite. Let $xy=1$ for some $x,\,y\in R$. As $yx$ is an idempotent, so for every $n\in\mathbb{N}$ we have $R(yx)^{n}=Ryx$ is an ideal of $R$. In particular, $yxy=y\in Ryx$, implying that $R=Ry\subseteq Ryx\subseteq Rx$. Thus $Rx=R$ and so $R$ is Dedekind finite. ∎ Further, if $R$ is an abelian ring, then the converse of Theorem 2.7 is also true.
@@
Since $a$ is strongly regular, we have $Ra=Re$ for some idempotent $e\in R$. As $R$ is an abelian ring, $Ra$ is an ideal. Hence, for any $r\in R$, $a^{n}r\in Ra^{n}$ for any $n\in\mathbb{N}$. ∎ However, it is not necessary that for a strongly regular element $a\in R$, $Ra^{n}$ is an ideal for some $n\in\mathbb{N}$. For example, if $R=\mathbb{M}_{2}(\mathbb{Q})$. Then $a=E_{11}$ is a strongly regular element in $R$, but $Ra^{n}=Ra$ is not an ideal for any $n\in\mathbb{N}.$
@@
Any ring with identity satisfying the ascending chain condition on left annihilators of subsets of $R$ is left Dischinger. In particular, left Noetherian rings are left Dischinger.
It is natural to ask whether a ring satisfying the ascending chain condition on left annihilators of subsets is right Dischinger. In Section 3, we will show that this may not be. Moreover, we prove that with some additional condition, a ring satisfying the condition in Theorem 2.11 becomes right Dischinger (see Corollary 2.14). For this, we need the following lemma. We will tacitly use the fact that if $ab$ is nilpotent, then so is $ba$ for any $a,\,b\in R$.
@@
Let $a$ be an element in a ring $R$ and $\rm{N}$$^{r}_{R}$$(a)$ be the subset of nilpotent elements such that for every $r\in\rm{N}$$^{r}_{R}$$(a)$ there exist elements $r_{i}\in R$, $i\in\mathbb{N}$ such that $r=r_{1}a$, $ar_{i}=r_{i+1}a$ and $ar_{k}=0$ for some $k\in\mathbb{N}$. Then $\rm{N}$$^{r}_{R}$$(a)$ is a nil subring of $R$.
It is easy to see that $\rm{N}$$^{r}_{R}$$(a)$ is closed under addition. Let $r,\,s\in\rm{N}$$^{r}_{R}$$(a)$. Then there exist $r_{i},\,s_{i}\in R$ for $i\in\mathbb{N}$ such that $r=r_{1}a$, $ar_{i}=r_{i+1}a$, $s=s_{1}a$, and $as_{i}=s_{i+1}a$ for all $i\in\mathbb{N}$. Then $ar_{i}as_{i}=r_{i+1}as_{i+1}a$ for all $i$. Let $k$ be the least positive integer such that one of the $ar_{k}$ or $as_{k}$ is zero. Then $ar_{k}as_{k}=0$ implying that $r_{1}as_{1}a\in\rm{N}$$^{r}_{R}$$(a)$. ∎ Analogously, we can define the left version of $\rm{N}$$^{r}_{R}$(a)(a)for any element $a\in R$ and denote it by $\rm{N}$$^{\ell}_{R}$$(a)$
@@
A ring $R$ with identity is called left Goldie if it satisfies the ascending chain condition on left annihilators and $_{R}R$ has finite uniform dimension. In [10], Lanski proved that every nil subring of a left Goldie ring with identity is nilpotent. In view of this and Theorem 2.13, we have the following result.
@@
A ring $R$ in which every nil subring is nilpotent is Dischinger. In particular, a left Goldie ring with identity is a Dischinger.
As a left Noetherian ring is left Goldie, so one-sided Noetherian rings with identity are Dischinger.